[Hallc_running] [New Logentry] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Wed Nov 21 14:00:01 EST 2018

Logentry Text:
Hello from the incoming RC - thank you to Eric for his RC duties and smooth transition.

Summary since yesterday:
- Beam since swing yesterday had several downtimes due to RF issues (see strip chart https://logbooks.jlab.org/entry/3629435) - e.g. swing had 223min ABU, 250min BNA and owl had 152min ABU, 319min BNA.
- MCC put beam off at ~22:45 for an extended time (1L24 power supply repair, 1L17 tuner repairs, MQC2A16 mismatch issue). See crew chief summary https://logbooks.jlab.org/entry/3629150 and https://logbooks.jlab.org/entry/3629153.
- Owl started with beam off and experienced several downtimes, including no beam from 5:50am until end of shift due to 1L17-8 issue (https://logbooks.jlab.org/entry/3629350).
- At 4.01am a Hall C Beamway HD Fire Alarm was triggered, a controlled access was made by experts. It was a false alarm and everything was reset (https://logbooks.jlab.org/entry/3629247).
- Day shift started with beam off due to the 1L17-8 issue and we had 45min downtime starting at ~10am caused by ARC4A tripping off (https://logbooks.jlab.org/entry/3629367).

- Data taking for kinematic group 15 has been on-going since end of day shift yesterday.
- We just finished kinematic group 15 and will change configuration to group 11 imminently and start running with this group.

Further notices:
- Hall A high beam current test happened this morning and is completed.
- Steve added EPICS signals to alarm handler to avoid a repeat occurrence of the previous missing alarm when SHMS Q2 tripped (https://logbooks.jlab.org/entry/3629011).

RC Meeting:
- We will have an RC meeting today at 4:30pm, and no further meetings until Monday.
- Until Mon we plan to have an informal meeting once per day in counting hut with spokespersons, unless there are major incidents or any special requests for a meeting. In which case please let me know.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser. The current plan is for all Halls to go to restricted access on Monday morning and for Halls D/C to be locked up again Tuesday Swing.

Planned accesses to the hall:
Monday, Nov 26: Work in the hall should be coordinated with Walter K. Aerogel change will take place (~8 hours duration Mon). HGC work will also take place, coordinating around aerogel change plans.
Tuesday, Nov 27: installing brushes near the spectrometer wheels for S. Lassiter (originally planned for Tue/Wed).
Please let me know if there are any other plans for work in the hall.

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463)
Vijay Kumar (Regina group) has requested a walkthrough during the next access


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3629460
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181121/80836bca/attachment.html>

More information about the Hallc_running mailing list